OpenStudy (anonymous):

HELLLPPPPP!!!!!!!!!! MEEEEE!!!!!!!!!!! Write an equation with a varable on both sides of the equal sign that has infintely many solutions. Solve the equation and explain why it has an infinite number of solutions.

OpenStudy (anonymous):

Medal to whoever helps

OpenStudy (anonymous):

x=x where x equal to any value so infinite values can be used

OpenStudy (anonymous):

can u give an example?

OpenStudy (anonymous):

If x=x you are saying that any value you give x equals itself. Since x is a variable that can be assigned any value, its value is infinite. examples are x= car so car=car; true x= moon so moon=moon ; also true or x equals infinity so infinity=infinity; true

OpenStudy (anonymous):

Better yet...x+3=x+3 is the same just like above.

OpenStudy (anonymous):

and if you want to take the question literally 3+x=x+3 where x is a constant that ranges from (-infinity, infinity)